OpenStudy (anonymous):

A flashlight has 6 batteries, 2 of which are defective. If 2 are selected at random without replacement, find the probability that both are defective.

OpenStudy (anonymous):

\[\frac{2}{6}\times \frac{1}{5}\]

OpenStudy (anonymous):

1/15

OpenStudy (anonymous):

i get \[\frac{1}{15}\]

OpenStudy (anonymous):

probability both are defective is probability first one is times probability second one is GIVEN that the first one is

OpenStudy (anonymous):

\begin{array}l\color{red}{\normalsize\text{r}}\color{orange}{\normalsize\text{i}}\color{#9c9a2e}{\normalsize\text{g}}\color{green}{\normalsize\text{h}}\color{blue}{\normalsize\text{t}}\color{purple}{\normalsize\text{?}}\color{purple}{\normalsize\text{}}\end{array}

OpenStudy (anonymous):

so its 1/15

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thanks